OpenStudy (141823):

If the first term is 10, the second is 1,000, the third is 100,000 and the fourth is 10,000,000. What is the 80th term.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

1st term= 1 zero 2nd term= 3 zeros 3rd term= 5zeros . . .

OpenStudy (anonymous):

nth term= a+(n-1)d zeros

OpenStudy (anonymous):

where n=80 d=2 a=1 80th term= 1+(79)2 = 159 zeros

OpenStudy (anonymous):

so the answer is 10^159

OpenStudy (141823):

Thank you, I understand it now!